ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60, 62 Bens̆ić, Mirta, Faculty of Economics, Osijek, Croatia Asymptotic properties of the least square estimators in regression models with singular errors There is now considerable theoretical and practical in analyzing regression models with long covariance stationary errors which posses nonparametric spectrum unbounded at some frequencies. Conditions for consistency of LSE in continuous time linear regression models with long-memory stationary Gaussian and non-Gaussian errors will be presented as well as a spectral representation of asymptotic distributions of the LSE. These results are joint work with Prof. N. N. Leonenko (University of Kiev, Ukraine). Reference: N. N. Leonenko, M. Bens̆ić, On estimation of regression coefficients of long memory random fields observed on the arrays, Random Oper. and Stoch. Equ., 5, 3 (1997), 237 - 252.